Jerry is experimenting with chemicals in the laboratory. He mixes a solution that is 10% acid with a solution that is 30% acid. How much of the 30% acid solution will be needed to make 40 liters of solution that is 25% acid?
Answer by stanbon(75887)

He mixes a solution that is 10% acid with a solution that is 30% acid. How much of the 30% acid solution will be needed to make 40 liters of solution that is 25% acid?
-----
Equation:
acid + acid = acid
-----
0.10x + 0.30(40-x) = 0.25*40
---
Multiply thru by 100 to get:
10x + 30*40 - 30x = 25*40
-----
-20x = -5*40
x = 10 liters (amt. of 10% solution needed)
---
40-x = 30 liters (amt. of 30^ acid needed)
============================================
Cheers,
Stan H.
============